OpenStudy (anonymous):

3x + 2y= 2 4x + y= -4

OpenStudy (anonymous):

multiply second equation by 2

OpenStudy (anonymous):

3x+2y=2 8x+2y=-8

OpenStudy (anonymous):

subtract second from first

OpenStudy (anonymous):

-5x=10 x=-2

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

um you did that wrong

OpenStudy (anonymous):

let me show you the correct way its called elimination method

OpenStudy (anonymous):

oh where's my mistake? i'm not perfect sorry.

OpenStudy (anonymous):

3x + 2y= 2 (4x + y= -4)-2

OpenStudy (anonymous):

3x + 2y= 2 4x + y= -4 --------- y = -4x - 4 3x + 2(-4x - 4) = 2 3x -8x - 8 = 2 -5x = 10 x = -2 3(-2) + 2y = 2 -6 + 2y = 2 2y = 8 y = 4 (-2,4)

OpenStudy (anonymous):

3x+2y=2 -8x-2y=8

OpenStudy (anonymous):

-5x=10 x=-2

OpenStudy (anonymous):

substitution >>> elimination

OpenStudy (anonymous):

you want the y's to cancel

OpenStudy (anonymous):

so its times -2 not 2

OpenStudy (anonymous):

are you not seeing how we got the same answer. Chris?

OpenStudy (anonymous):

the difference is I subtracted the equations. I didn't add them.

OpenStudy (anonymous):

slet me show you substitution then

OpenStudy (anonymous):

You can do either. You might want to properly educate yourself on the topic before you start calling people out.

OpenStudy (anonymous):

3x + 2y= 2 4x + y= -4

OpenStudy (anonymous):

wait5 a sec

OpenStudy (anonymous):

ill show you

OpenStudy (anonymous):

i already did the substitution method juss sayin'

OpenStudy (anonymous):

for substitution i would solve one equation for y

OpenStudy (anonymous):

hah heated.

OpenStudy (anonymous):

before you finish. my answer was x = -2 and y = 4 as shown above. So you are going to tell me that is incorrect?

OpenStudy (anonymous):

4x + y= -4 y=-4x-4

OpenStudy (anonymous):

substitute that for y in the other equation

OpenStudy (anonymous):

4 = -4x-4 8 = -4x x=-2

OpenStudy (anonymous):

Chris. just stop.

OpenStudy (anonymous):

3x + 2(-4x-4)= 2 3x-8x-8=2 -5x=10 x=-2

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

thank you for again proving my point

OpenStudy (anonymous):

thats substitution

OpenStudy (anonymous):

there was nothing wrong with the elimination method I did. get over it.

OpenStudy (anonymous):

you multiplies by 2 instead of -2

OpenStudy (anonymous):

it was completely wrong

OpenStudy (anonymous):

+2y and +2y do not eliminate

OpenStudy (anonymous):

that would give him 4y

OpenStudy (anonymous):

I am not saying you were trying to do it wrong, you just made a mistake

OpenStudy (radar):

ChrisV you can always expand your horizons. Yes you can multiply by -2 then add, But you can just as easily and correctly multiply by +2 and then subtract. The fact that you can eliminate by both adding or subtracting is something you should put in your "solving" tool box.

OpenStudy (anonymous):

3x+2y=2 8x+2y=-8 11x+4y=-6 that in know way is right

OpenStudy (anonymous):

you can not multiply by 2 and get the right answer there

OpenStudy (anonymous):

if you can my friend you have invented a new form of the elimination method

OpenStudy (anonymous):

please add it to my precal book so i can inform the aurhors

OpenStudy (anonymous):

you have to multiply bu -2 not 2

OpenStudy (anonymous):

3x+2y=2 -8x-2y=8 -5x=10 x=-2

OpenStudy (radar):

You are missing the point: 3x+2y=2 -(8x+2y=-8) ------------ -5x = 10 x=-2

OpenStudy (anonymous):

why not call it the subtraction method then? The point is to multiply by a value that eliminates a variable

OpenStudy (anonymous):

Just tryinying to sho the simplest way i guess, but yes that comes out correct too

OpenStudy (anonymous):

show

OpenStudy (anonymous):

anyways lets do something else lol, nothing more to argue about here. Just gonna disagree anyways

OpenStudy (radar):

Yes that is true, and you may eliminate it either by subtracting or adding depending on your choice of the sign of your multiplier. I am not saying you are wrong and if you wish to restrict your mode of operation to always adding, then have at it.

OpenStudy (anonymous):

not adding its called multiplication

OpenStudy (anonymous):

lol Chris, maybe when you further your actual understanding of what it means to execute the elimination method then talk to me. Go ask any knowledgeable math teacher or professor you have ever had. I don't appreciate being told that I'm wrong when clearly, I am not. I'm not at all saying what you did was wrong but the fact that you have the inability to realize that the elimination method has multiple execution techniques and then accusing EVERYONE else of being wrong is just disrespectful. I'm sure you are a smart guy but you need to expand your horizon. I have been doing a lot of math for a very long time now. Good luck in the future.